OpenStudy (anonymous):

Solve for x in the proportion x over the quantity x plus 12 equals 1 over the quantity x minus 1. A:x ≈ −6.02 and x ≈ 5.02 B:x ≈ −3.46 and x ≈ 3.46 C:x ≈ −3.14 and x ≈ 4.14 D:x ≈ −2.61 and x ≈ 4.61

OpenStudy (campbell_st):

\[\frac{x}{x + 12} = \frac{1}{x - 1}\] cross multiply to remove the denominators. \[x( x-1) = 1(x + 12)\] expand and simplify \[x^2 - x=x + 12\] rewrite the equation will all terms on the left \[x^2 - 2x - 12 = 0\] you will need to solve this using the general quadratic formula \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] you need to substitute a = 1, b = -2 and c = -12 and solve to get you answers

OpenStudy (anonymous):

Ok,Thankss(: